# The Open Journal of Astrophysics

### v1

@@ Abstract
The correlations between the positions and shapes of galaxies, i.e. intrinsic alignments, have been measured in many observational studies and hydrodynamical simulations. The alignments of disk galaxies in hydrodynamical simulations have been measured to be positive, null and negative with varying methodologies, samples and hydrodynamical simulations. This work compares the correlations of disks and ellipticals around all galaxies and disks around ellipticals at $z=0$ and $z=1$ for simple and reduced shapes in TNG300, Horizon-AGN and EAGLE for multiple morphological definitions in a consistent way. All types of signals are positive and robust in TNG300 and EAGLE and positive or null in Horizon-AGN, except for the disks around ellipticals correlation for reduced shapes at $z=1$ when defined by $|v/\sigma|$, which is negative. A re-weighting of the ellipticals around galaxies signals in TNG300, according to the underlying stellar mass distributions of the samples, highlights the importance of the influence of (sub-grid) physics at these non-linear scales.
@@ 1 Introduction
One way to gain understanding of galaxy evolution is through the measurement of IA in hydrodynamical simulations, which allow us to study galaxies in ways that observations do not (yet) allow. For nearly a decade, IA has been studied in hydrodynamical simulations to gain insight on the correlation between stellar mass, morphology, luminosity, AGN feedback, among other factors, and the amplitude of IA (e.g Tenneti et al., 2014; Chisari et al., 2015; Shao et al., 2016; Tenneti et al., 2017; Chisari et al., 2018; Shi et al., 2021). In general, the IA amplitude is always highest for elliptical, high mass and luminous galaxies (e.g. Chisari et al., 2016; Tenneti et al., 2016; Samuroff et al., 2021). Furthermore, the redshift evolution of the IA correlation functions has also been studied, which is crucial for its mitigation in weak lensing studies (e.g. Welker et al., 2017; Bhowmick et al., 2019; Bate et al., 2019).
@@
One of the open problems in this field is the alignment of disks, which might be the majority of weak lensing sources for Stage IV surveys. Disk alignments have been measured by various studies in differing simulations and for a variety of morphological definitions and methodologies, with conflicting positive (radial), negative (tangential) and null detections. Tenneti et al. (2016) found positive disk around elliptical ellipticity-direction alignments in MassiveBlack-II (Khandai et al., 2015) and Illustris (Genel et al., 2014; Vogelsberger et al., 2014), when splitting galaxies via their bulge-to-total ratio. Hilbert et al. (2017) also found a positive disk alignment in Illustris, mimicking observations and selecting disks by their BPZ spectral type. The alignment amplitude obtained from modelling the correlation functions with the nonlinear alignment model (NLA) at $z>0.5$ yield positive alignments in Illustris, MassiveBlack-II and TNG300 (Nelson et al., 2021) for blue (disk) galaxies in the work of Samuroff et al. (2021). When measuring the 3D power spectrum for galaxy alignments, Shi et al. (2021) find a disk alignment consistent with zero in TNG300, when defining disks using $\kappa_{\mathrm{rot}}$ (see Eq. (13)). They do find a ’spin flip’ signal, where the spin, i.e. the direction of the angular momentum, of low-mass disks is parallel to the filament and high mass disks are aligned at $z=0.3,0.5$. In the newest TNG simulation, MTNG740, Delgado et al. (2023) measure a positive disk IA signal for disks defined by $\kappa_{\mathrm{rot}}$ and sSFR. Chisari et al. (2015) found a negative disk alignment when measuring the correlations of disks around ellipticals, defined by $v/\sigma$ (see Sec. 2.4 for measurement method), in Horizon-AGN (Dubois et al., 2014) for reduced shapes, which are also found at higher redshift in Chisari et al. (2016). Dubois et al. (2014) measures the alignment of galaxies in Horizon-AGN and its dependence on various properties, e.g. $v/\sigma$, sSFR and colour, finding a positive spin-filament alignment for disks. A positive alignment between the spin of blue galaxies and the filament is also found by Welker et al. (2017) in Horizon-AGN, who also looked extensively into satellite alignments. Kraljic et al. (2020) also measures the alignment between galaxy spin and the filaments of the matter distribution and finds that in SIMBA (Davé et al., 2019), the alignment signal flips between high and low stellar mass samples, for both high and low $v/\sigma$ populations of galaxies. In EAGLE (Schaye et al., 2015; Crain et al., 2015), Velliscig et al. (2015) measured the 3D alignment of galaxies based on their axis ratios and found a positive disk signal.
@@
In observations, most disk alignments are found to be consistent with zero. Using the SDSS Main sample, Mandelbaum et al. (2006) measured no significant blue alignment for blue galaxies around the full sample, but did find the red and blue alignments consistent in the highest luminosity bin, which is attributed to the fact that these galaxies are almost red. Later, Mandelbaum et al. (2011) also found no alignment signal for selection of WiggleZ Dark Energy Survey galaxies that overlaps with SDSS, constraining the blue alignment in a model-dependent way. Johnston, Harry et al. (2019) performed the measurement of intrinsic alignments of blue galaxies in KiDS+GAMA, leading to tighter constraints of the lack of alignment of blue galaxies. Recently, Georgiou et al. (2025), who measured the disk signal in the KiDS-1000 bright sample for both blue and $n_{s}<2.5$ galaxies, also found no significant alignment for this population of galaxies.
@@
This work is structured as follows. Section 2 focusses on the simulation data (Sec. 2.1); measurement methodologies of galaxy shapes (Sec. 2.2); projected correlation functions (Sec. 2.3) and the variables used in the morphological definitions, which are used to select the samples (Sec. 2.4). Section 3 presents the results of the distributions of the morphological variables (Sec. 3.1); measurement of IA correlations of disks and ellipticals around galaxies (Sec. 3.2); disks around ellipticals (Sec. 3.3); the disk signal in Horizon-AGN (Sec. 3.4) and the influence of the stellar mass distributions on the ellipticals around galaxies measurements (Sec. 3.5). Section 4 provides discussion and an outlook for future research and the conclusions are summarised in Section 5.
@@ 2.1 Simulations
Throughout this work, we use the data from three hydrodynamical simulation suites: IllustrisTNG (Nelson et al., 2021) and EAGLE (Schaye et al., 2015; Crain et al., 2015), which are publicly available, and Horizon-AGN (Dubois et al., 2014). These simulations are run using a variety of parameters, codes and models, which are described briefly below. For more details, see the cited papers.
@@ 2.1.1 IllustrisTNG
The IllustrisTNG project contains simulations with multiple boxsizes and resolutions, of which TNG300-1 is used throughout this work, and TNG100-1 and TNG100-2 are used in Appendix A to gauge the influence of boxsize and resolution effects on the results. The volume of TNG300 is a box with length $205\mathrm{cMpc}/h$ on each side, containing $2500^{3}$ dark matter particles of mass $m_{\mathrm{DM}}=5.9\times 10^{7}\mathrm{M_{\odot}}$, which are evolved using the AREPO code (Springel, 2010) from $z=127$ to $z=0$. The cosmological parameters are assumed to be: $\Omega_{\mathrm{m},0}=0.3089$, $\Omega_{\mathrm{b},0}=0.0486$, $\sigma_{8}=0.8159$, $n_{\mathrm{s}}=0.9667$ and $h=0.6774$ (Planck Collaboration et al., 2016). Physical processes that are not resolved are governed by sub-grid models, which are described in Pillepich et al. (2017) and Weinberger et al. (2016) and include procedures for radiative cooling and heating; star formation; stellar evolution and feedback; chemical enrichment and black hole accretion, mergers and AGN feedback with thermal (quasar) and kinetic (wind) modes. Haloes, sub-haloes and galaxies are found and defined using the Friends-of-Friends (FoF) and SUBFIND algorithms (Springel et al., 2001), which group dark matter particles and then determine the gravitationally bound (sub-)structures.
@@ 2.1.3 Horizon-AGN
The Horizon-AGN simulation (Dubois et al., 2014) has a volume of $100^{3}[\mathrm{cMpc}/h]^{3}$ and $1024^{3}$ dark matter particles of mass $m_{\mathrm{DM}}=8\times 10^{7}\mathrm{M_{\odot}}$. Horizon-AGN uses the Wilkinson Microwave Anisotropy Probe year 7 (Komatsu et al., 2011, WMAP7) cosmological parameters: $\Omega_{\mathrm{m}}=0.272$, $\Omega_{\mathrm{b}}=0.045$, $\sigma_{8}=0.81$, $n_{\mathrm{s}}=0.967$ and $h=0.704$. This simulation is run using the adaptive mesh refinement code RAMSES (Teyssier, 2002), which evolves gas cells that are resolved further depending on the number of dark matter particles. Similar to the other two simulations, sub-grid models govern unresolved physics including gas cooling and heating; stellar winds; star formation; supernova feedback; black hole accretion and AGN feedback in kinematic ‘radio’ and thermal ‘quasar’ modes. Galaxies are found using the AdaptaHOP algorithm (Aubert et al., 2004), which uses the distribution of stellar particles to identify galaxies with a minimum of 50 stellar particles.
@@ 2.2 Shapes of galaxies
For consistency, the shapes used in this work have all been calculated using the same method for each simulation. In this entire work, the projected shapes are calculated using the simple (Eq. (1)) or reduced inertia tensor (Eq. (2)), defined by:
$I^{k,l}_{j}$ $=\frac{1}{M_{j}}\Sigma^{N_{j}}_{i=1}m_{i}r^{k}_{i}r^{l}_{i}$ (1)
$I^{k,l}_{j}$ $=\frac{1}{M_{j}}\Sigma^{N_{j}}_{i=1}m_{i}\frac{r^{k}_{i}r^{l}_{i}}{r^{2}}$ (2)
where $k,l$ run from 0 to $1$ and denote the $x,y$ (or $z$) components of vector $\mathbf{r_{i}}$, depending on the projection axis (line-of-sight), which is chosen to be $z$ in this work. $\mathbf{r_{i}}$ is the position of each particle $i$ in galaxy $j$ relative to the centre of mass of the galaxy ($\mathbf{r_{i}}=\mathbf{x_{i}}-\mathbf{x_{j}}$), with $i$ running from $1$ to $N_{j}$. The mass-weighted centre of mass of the stellar particles for each galaxy ($j$) is calculated as follows:
$\mathbf{x}_{j}=\frac{1}{M_{j}}\Sigma^{N_{j}}_{i=1}m_{i}\mathbf{x}_{i}$ (3)
where $M_{j}$ is the galaxy stellar mass ($\sum_{i}m_{i}$) and $\mathbf{x_{i}}$ and $m_{i}$ are the stellar particle positions and masses. In TNG300, the ‘wind-particles’, which are actually gas particles, have been omitted from the stellar mass.
The eigenvectors of the simple (reduced) inertia tensor, Equation (1) ((2)), are the unit vectors that denote the directions of the semi-major and semi-minor axes in the galaxy shape. For the simple inertia tensor, the axis lengths are given by the square root of the eigenvalues and named, $a$ (semi-major axis length) and $b$ (semi-minor).
@@ 2.3 Projected correlation functions
To match what is measured in observations, we use projected statistics of galaxy alignments. In this work, we use two statistics to quantify the alignments of galaxy shapes around galaxy positions: $\tilde{\xi}_{g+,2}$ (the quadrupole) and $w_{g+}$ (only in Appendix B). The correlation functions are estimated using the Landy-Szalay estimators (Landy & Szalay, 1993), which use a shape sample, $S$, a position sample, $D$ and a random sample, $R$, to estimate the correlation of galaxy shapes and positions, separated by $\bf{r}$. Both statistics use the same estimator, but are binned in different ways: either in $(r_{p},\Pi)$, where $r_{p}$ is the length of $\bf{r}$, projected along the line-of-sight, $\Pi$, or in $(r,\mu_{r})$, where $r$ is the length of $\bf{r}$ and $\mu_{r}=\Pi/r$, the cosine of the angle between $\bf{r}$ and $\Pi$. Regardless of the binning, the modified Landy-Szalay estimator used in this work (Mandelbaum et al., 2011) is defined as:
@@
Here, $(r,\mu_{r})$ can be substituted for $(r_{p},\Pi)$. As shown in Equations (4) - (6), the correlation functions can be estimated by first finding all galaxy pairs $\{i,j\}$ between the position and shape sample in an $(r_{p},\Pi)$ bin. For each pair, the tangential/radial component of the ellipticity of the shape galaxy $j$ (given position $i$) can be obtained using Equation (6), where $q$ is the axis ratio of the projected shape, $q=b/a$, as defined in Section 2.2 and $\phi$ is the angle between the separation vector $\bf{r}$ and the semi-major axis direction (also defined in Section 2.2). A summation of $e_{+}$ over all galaxy pairs within an $(r_{p},\Pi)$ or $(r,\mu_{r})$ bin, is then divided by the responsivity factor $\mathcal{R}$, defined as $\mathcal{R}=1-\langle e^{2}\rangle$, when unweighted, with $e=\frac{1-q^{2}}{1+q^{2}}$, to obtain $S_{+}D$ (Eq. (5)). Each galaxy in the shape and position sample has weight $w_{j},w_{i}$, respectively, which are set to $1$ in all sections, except in Section 3.5. If the weights are not unity, $\mathcal{R}$ is defined as $\mathcal{R}=\frac{\sum_{j}{w_{j}(1-e^{2}/2)}}{\sum_{j}{w_{j}}}$, in line with Jarvis et al. (2003). Equivalently to $S_{+}D$, $S_{+}R$ is the summation over galaxy ellipticities around random points, which is be assumed to be zero in this work, following Chisari et al. (2015).

### v2

@@ Abstract
The correlations between the positions and shapes of galaxies, i.e. intrinsic alignments, have been measured in many observational studies and hydrodynamical simulations. The position-shape correlation measurements of disk galaxies with varying methodologies, samples and hydrodynamical simulations are inconsistent in amplitude and sign. This work compares the correlations of disk and elliptical shapes around all galaxy positions and disk shapes around the positions of ellipticals at $z=0$ and $z=1$ for two different shape definitions in TNG300, Horizon-AGN and EAGLE for multiple morphological definitions in a consistent way. All types of signals are positive and robust in TNG300 and EAGLE and positive or null in Horizon-AGN. The exception to this is the negative correlation of disks around ellipticals in Horizon-AGN. This arises for reduced shapes, which down-weight the outskirts of galaxies, at $z=1$, when disks are identified via a threshold in $|v/\sigma|$, the rotational velocity over the velocity dispersion. A re-weighting of the ellipticals around all galaxies signals in TNG300, according to the underlying stellar mass distributions of the samples, highlights the importance of the influence of (sub-grid) physics at these non-linear scales.
@@ 1 Introduction
One way to gain understanding of galaxy evolution is through the measurement of IA in hydrodynamical simulations, which allow us to study galaxies in ways that observations do not (yet) allow. For a decade, IA has been studied in hydrodynamical simulations to gain insight on the correlation between stellar mass, morphology, luminosity, AGN feedback, among other factors, and the amplitude of IA (e.g Tenneti et al., 2014; Chisari et al., 2015; Shao et al., 2016; Tenneti et al., 2017; Chisari et al., 2018; Shi et al., 2021). In general, the IA amplitude is always highest for elliptical, high mass and luminous galaxies (e.g. Chisari et al., 2016; Tenneti et al., 2016; Samuroff et al., 2021). Furthermore, the redshift evolution of the IA correlation functions has also been studied, which is crucial for its mitigation in weak lensing studies (e.g. Welker et al., 2017; Bhowmick et al., 2020; Bate et al., 2020).
@@
One of the open problems in this field is the alignment of disks, which might be the majority of weak lensing sources for Stage IV surveys. Disk alignments have been measured by various studies in differing simulations and for a variety of morphological definitions and methodologies, with conflicting positive (radial), negative (tangential) and null detections. Tenneti et al. (2016) found positive disk around elliptical ellipticity-direction alignments in MassiveBlack-II (Khandai et al., 2015) and Illustris (Genel et al., 2014; Vogelsberger et al., 2014), when splitting galaxies via their bulge-to-total ratio. Hilbert et al. (2017) also found a positive disk alignment in Illustris, mimicking observations and selecting disks by their Bayesian Photometric redshift (BPZ) spectral type. The alignment amplitude obtained from modelling the correlation functions with the nonlinear alignment model (NLA) at $z>0.5$ yield positive alignments in Illustris, MassiveBlack-II and TNG300 (Nelson et al., 2019) for blue (disk) galaxies in the work of Samuroff et al. (2021). When measuring the 3D power spectrum for galaxy alignments, Shi et al. (2021) find a disk alignment consistent with zero in TNG300, when defining disks using $\kappa_{\mathrm{rot}}$ (see Eq. 13). They do find a ’spin flip’ signal, where the spin, i.e. the direction of the angular momentum, of low-mass disks is parallel to the filament and high mass disks are aligned at $z=0.3,0.5$. In the newest TNG simulation, MTNG740, Delgado et al. (2023) measure a positive disk IA signal for disks defined by $\kappa_{\mathrm{rot}}$ and sSFR. Chisari et al. (2015) found a negative disk alignment when measuring the correlations of disks around ellipticals, defined by $v/\sigma$ (see Sec. 2.4 for measurement method), in Horizon-AGN (Dubois et al., 2014) for reduced shapes, which are also found at higher redshift in Chisari et al. (2016). Dubois et al. (2014) measures the alignment of galaxies in Horizon-AGN and its dependence on various properties, e.g. $v/\sigma$, sSFR and colour, finding a positive spin-filament alignment for disks. Kraljic et al. (2020) measures the alignment between galaxy spin and the filaments of the matter distribution and finds that in SIMBA (Davé et al., 2019), the alignment signal flips between high and low stellar mass samples, for both high and low $v/\sigma$ populations of galaxies. A positive alignment between the spin of blue galaxies and the filament is also found by Welker et al. (2017) in Horizon-AGN, who also looked extensively into satellite alignments. In EAGLE (Schaye et al., 2015; Crain et al., 2015), Velliscig et al. (2015) measured the 3D alignment of galaxies based on their axis ratios and found a positive disk signal.
@@
In observations, most disk alignments are found to be consistent with zero. Using the SDSS Main sample, Mandelbaum et al. (2006) measured no significant blue alignment for blue galaxies around the full sample, but did find the red and blue alignments consistent in the highest luminosity bin, which is attributed to the fact that these galaxies are almost red. Later, Mandelbaum et al. (2011) also found no alignment signal for selection of WiggleZ Dark Energy Survey galaxies that overlaps with SDSS, constraining the blue alignment in a model-dependent way. Johnston et al. (2019) performed the measurement of intrinsic alignments of blue galaxies in KiDS+GAMA, leading to tighter constraints of the lack of alignment of blue galaxies. Recently, Georgiou et al. (2025), who measured the disk signal in the KiDS-1000 bright sample for both blue and $n_{s}<2.5$ galaxies, also found no significant alignment for this population of galaxies.
@@
This work is structured as follows. Section 2 focusses on the simulation data (Sec. 2.1); measurement methodologies of galaxy shapes (Sec. 2.2); projected correlation functions (Sec. 2.3) and the variables used in the morphological definitions, which are used to select the samples (Sec. 2.4). Section 3 presents the results of the distributions of the morphological variables (Sec. 3.1); measurement of IA correlations of disks and ellipticals around all galaxies (Sec. 3.2); disks around ellipticals (Sec. 3.3); the disk signal in Horizon-AGN (Sec. 3.4) and the influence of the stellar mass distributions on the ellipticals around all galaxies measurements (Sec. 3.5). Section 4 provides discussion and an outlook for future research and the conclusions are summarised in Section 5.
@@ 2.1 Simulations
Throughout this work, we use the data from three hydrodynamical simulation suites: IllustrisTNG (Nelson et al., 2019) and EAGLE (Schaye et al., 2015; Crain et al., 2015), which are publicly available, and Horizon-AGN (Dubois et al., 2014). We chose these simulations because they are most widely used when studying intrinsic alignments, allowing for comparison to previous works. Their resolutions and boxsizes allow for large enough samples of a representative group of galaxies, i.e. not just massive ones, but milky-way-like as well, with a enough signal-to-noise for a robust detection of intrinsic alignment correlation functions at non-linear scales. These simulations are run using a variety of parameters, codes and models, which are described briefly below and summarised in Table 1. For more details, see the cited papers.
Table 1: Parameter information of each simulation used in this work, including cosmology, boxsize, number of dark matter (DM) particles, DM particle mass, numerical scheme, AGN feedback model modes and level of calibration.
Simulation Cosmology Boxsize Number of DM particle Numerical scheme AGN feedback Calibration
DM particles mass level
TNG300 Planck 2016 205 $\mathrm{cMpc}/h$ $2500^{3}$ $5.9\times 10^{7}\mathrm{M_{\odot}}$ moving mesh thermal and kinetic high
TNG100-1 Planck 2016 75 $\mathrm{cMpc}/h$ $1820^{3}$ $7.5\times 10^{6}\mathrm{M_{\odot}}$ moving mesh thermal and kinetic high
TNG100-2 Planck 2016 75 $\mathrm{cMpc}/h$ $910^{3}$ $6.0\times 10^{7}\mathrm{M_{\odot}}$ moving mesh thermal and kinetic high
EAGLE Planck 2014 67.77 $\mathrm{cMpc}/h$ $1504^{3}$ $6.6\times 10^{6}\mathrm{M_{\odot}}$ smoothed particle hydrodynamics thermal low
Horizon-AGN WMAP7 100 $\mathrm{cMpc}/h$ $1024^{3}$ $8\times 10^{7}\mathrm{M_{\odot}}$ adaptive mesh refinement thermal and kinetic low
@@ 2.1.1 IllustrisTNG
The IllustrisTNG project contains simulations with multiple boxsizes and resolutions, of which TNG300-1 is used throughout this work, and TNG100-1 and TNG100-2 are used in Appendix A to gauge the influence of boxsize and resolution effects on the results. The volume of TNG300 is a box with length $205\mathrm{cMpc}/h$ on each side, containing $2500^{3}$ dark matter particles of mass $m_{\mathrm{DM}}=5.9\times 10^{7}\mathrm{M_{\odot}}$, which are evolved using the AREPO code (Springel, 2011) from $z=127$ to $z=0$. The cosmological parameters are assumed to be: $\Omega_{\mathrm{m},0}=0.3089$, $\Omega_{\mathrm{b},0}=0.0486$, $\sigma_{8}=0.8159$, $n_{\mathrm{s}}=0.9667$ and $h=0.6774$ (Planck Collaboration et al., 2016). Physical processes that are not resolved are governed by sub-grid models, which are described in Pillepich et al. (2018) and Weinberger et al. (2017) and include procedures for radiative cooling and heating; star formation; stellar evolution and feedback; chemical enrichment and black hole accretion, mergers and AGN feedback with thermal (quasar) and kinetic (wind) modes. Haloes, sub-haloes and galaxies are found and defined using the Friends-of-Friends (FoF) and SUBFIND algorithms (Springel et al., 2001), which group dark matter particles and then determine the gravitationally bound (sub-)structures. The IllustrisTNG simulations have been calibrated to the $z=0$ galaxy stellar mass function; stellar mass - halo mass relation; stellar mass - stellar size relation; stellar mass - BH mass relation and total gas mass content (in $r_{500}$) in massive galaxy groups and the shape of the cosmic star formation rate density at $z\leq 10$.
@@ 2.1.3 Horizon-AGN
The Horizon-AGN simulation (Dubois et al., 2014) has a volume of $100^{3}[\mathrm{cMpc}/h]^{3}$ and $1024^{3}$ dark matter particles of mass $m_{\mathrm{DM}}=8\times 10^{7}\mathrm{M_{\odot}}$. Horizon-AGN uses the Wilkinson Microwave Anisotropy Probe year 7 (Komatsu et al., 2011, WMAP7) cosmological parameters: $\Omega_{\mathrm{m}}=0.272$, $\Omega_{\mathrm{b}}=0.045$, $\sigma_{8}=0.81$, $n_{\mathrm{s}}=0.967$ and $h=0.704$. This simulation is run using the adaptive mesh refinement code RAMSES (Teyssier, 2002), which evolves gas cells that are resolved further depending on the number of dark matter particles. Similar to the other two simulations, sub-grid models govern unresolved physics including gas cooling and heating; stellar winds; star formation; supernova feedback; black hole accretion and AGN feedback in kinetic (radio) and thermal (quasar) modes. Galaxies are found using the AdaptaHOP algorithm (Aubert et al., 2004), which uses the distribution of stellar particles to identify galaxies with a minimum of 50 stellar particles. Horizon-AGN has been calibrated to reproduce the $M-\sigma$ relation at $z=0$.
@@ 2.2 Shapes of galaxies
For consistency, the shapes used in this work have all been calculated using the same method for each simulation. In this entire work, the projected shapes are calculated using the simple, Eq. 1, or reduced inertia tensor, Eq. 2, defined by:
$I^{k,l}_{s}$ $=\frac{1}{M_{s}}\sum^{N_{s}}_{n=1}m_{n}r^{k}_{n}r^{l}_{n}$ (1)
$\tilde{I}^{k,l}_{r}$ $=\frac{1}{M_{s}}\sum^{N_{s}}_{n=1}m_{n}\frac{r^{k}_{n}r^{l}_{n}}{r^{2}}$ (2)
where $k,l$ run from 0 to $1$ and denote the $x,y$ (or $z$) components of vector $\boldsymbol{r_{n}}$, depending on the projection axis (line-of-sight), which is chosen to be $z$ in this work. $\boldsymbol{r_{n}}$ is the position of each particle $n$ in galaxy $s$ relative to the centre of mass of the galaxy ($\boldsymbol{r_{n}}=\boldsymbol{x_{n}}-\boldsymbol{x_{s}}$), with $n$ running from $1$ to $N_{s}$. The mass-weighted centre of mass of the stellar particles for each galaxy ($s$) is calculated as follows:
$\boldsymbol{x}_{s}=\frac{1}{M_{s}}\sum^{N_{s}}_{n=1}m_{n}\boldsymbol{x}_{n}$ (3)
where $M_{s}$ is the galaxy stellar mass ($\sum_{n}m_{n}$) and $\boldsymbol{x_{n}}$ and $m_{n}$ are the stellar particle positions and masses. In TNG300, the ‘wind-particles’, which are actually gas particles, have been omitted from the stellar mass.
The eigenvectors of the simple (reduced) inertia tensor, Equation 1 (2), are the unit vectors that denote the directions of the semi-major and semi-minor axes in the galaxy shape. For the simple inertia tensor, the axis lengths are given by the square root of the eigenvalues and named, $a$ (semi-major axis length) and $b$ (semi-minor).
@@ 2.3 Projected correlation functions
To match what is measured in observations, we use projected statistics of galaxy alignments. In this work, we use two statistics to quantify the alignments of galaxy shapes around galaxy positions: $\tilde{\xi}_{g+,2}$ (the quadrupole) and $w_{g+}$ (only in Appendix B). The correlation functions are estimated using the Landy-Szalay estimators (Landy & Szalay, 1993), which use a shape sample, $S$, a position sample, $D$ and a random sample, $R$, to estimate the correlation of galaxy shapes and positions, separated by $\boldsymbol{r}$. Both statistics use the same estimator, but are binned in different ways: either in $(r_{p},\Pi)$, where $r_{p}$ is the length of $\boldsymbol{r}$, projected along the line-of-sight, $\Pi$, or in $(r,\mu_{r})$, where $r$ is the length of $\boldsymbol{r}$ and $\mu_{r}=\Pi/r$, the cosine of the angle between $\boldsymbol{r}$ and $\Pi$. Regardless of the binning, the modified Landy-Szalay estimator used in this work (Mandelbaum et al., 2011) is defined as:
@@
Here, $(r,\mu_{r})$ can be substituted for $(r_{p},\Pi)$. As shown in Equations 4 - 6, the correlation functions can be estimated by first finding all galaxy pairs $\{i,j\}$ between the position and shape sample in an $(r_{p},\Pi)$ bin. For each pair, the tangential/radial component of the ellipticity of the shape galaxy $j$ (given position $i$) can be obtained using Equation 6, where $q$ is the axis ratio of the projected shape, $q=b/a$, as defined in Section 2.2 and $\phi$ is the angle between the separation vector $\boldsymbol{r}$ and the semi-major axis direction (also defined in Section 2.2). A summation of $e_{+}$ over all galaxy pairs within an $(r_{p},\Pi)$ or $(r,\mu_{r})$ bin, is then divided by the responsivity factor $\mathcal{R}$, defined as $\mathcal{R}=1-\langle e^{2}\rangle$, when unweighted, with $e=\frac{1-q^{2}}{1+q^{2}}$, to obtain $S_{+}D$ (Eq. 5). Each galaxy in the shape and position sample has weight $w_{j},w_{i}$, respectively, which are set to $1$ in all sections, except in Section 3.5. If the weights are not unity, $\mathcal{R}$ is defined as $\mathcal{R}=\frac{\sum_{j}{w_{j}(1-e^{2}/2)}}{\sum_{j}{w_{j}}}$, in line with Jarvis et al. (2003). Equivalently to $S_{+}D$, $S_{+}R$ is the summation over galaxy ellipticities around random points. This term is be assumed to be zero in this work, following Chisari et al. (2015).